Ultrasonic motion sensor Pre - laboratory Assignment In experiments 1 and 2, you will use a cart, track, and motion detector(b) Assuming that there is no friction and that the objects velocity at the 0.75 m position is 0.75 m/s, what is the object
Answer #1

What I wook done by a Pre-laboratory assignment - theoemem > Change in ki Eo, ak all 1 7 using the woek-energy the forces onusing 2 kg - ki=w kg = k kitw [ki = 1/2 = 1/2 m. = 1X0.194 (0.75 +1.2 - 1.42 Andwa final kinetic eneegy encegy is 1.42 Joube